-1*5+1-6+15-30*-5/3
Answer by richwmiller(17219)

-1*5+1-6+15-30*-5/3
let's take it one step at a time
let's do the multiplications first
-1*5=-5
-30*-5/3
30/3=10
10*5=50
-*-=+
-30*-5/3=50
-5+1-6+15+50
combine
-5+1=-4
-4-6+15+50
-4-6=-10
-10+15+50
-10+15=5
5+50=55
so
-1*5+1-6+15-30*-5/3=55